# Towards a robust frequency-domain analysis: Spectral Rényi divergence revisited

### v1

@@ 3.1 Robustness of the trajectory of the optimization
Consider the gradient descent update with a sequence of learning rates $\gamma_{1},\gamma_{2},\ldots>0$ of $\widehat{\theta}_{1}$: for the $(m+1)$-th step,
@@
The proof is given in Appendix A. This instability remains for the second-order methods such as the Newton–Raphson method and the Fisher scoring (natural gradient) method.
Consider next the gradient descent update with a sequence of learning rates $\gamma_{1},\gamma_{2},\ldots>0$ of $\widehat{\theta}_{\alpha}$: for the $(m+1)$-th step,
with
$\mathcal{G}_{\alpha}(\theta\,;\,\widetilde{I}_{n})=\frac{\alpha}{(1-\alpha)n}\sum_{\omega\in\Omega_{n}}\left[1-\frac{S_{\theta}(\omega)}{\alpha S_{\theta}(\omega)+(1-\alpha)\widetilde{I}_{n}(\omega)}\right]\nabla_{\theta}\log S_{\theta}(\omega).$
The subsequent theorem implies that for any value of $\alpha\in(0,1)$, the path of the gradient descent for $\widehat{\theta}_{\alpha}$ is robust against the presence of outliers in the frequency domain. Let $\|\cdot\|_{\mathrm{op}}$ denote the operator norm.
Theorem 3.
Assume the following hold:
[…]
$\limsup_{n\to\infty}\sup_{\theta\in\Theta}\frac{1}{n}\sum_{\omega\in\Omega_{n}}\|\nabla^{2}_{\theta}\log S_{\theta}(\omega)\|_{\mathrm{op}}<\infty.$
Assume also for any $z>0$ and $\omega^{*}\in\Omega_{n}$, the gradient decent sequence $\{\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}^{z,\omega^{*}}]\}_{m=0,1,2,\ldots}$ lies in $\Theta$. Then, for any $m\in\mathbb{N}$, we have
$\lim_{n\to\infty}\sup_{z>0}\left\|\widehat{\theta}_{\alpha}^{(m)}[\widetilde{I}^{z,\omega^{*}}_{n}]-\widehat{\theta}_{\alpha}^{(m)}[\widetilde{I}_{n}]\right\|=0\text{ almost surely}.$ (6)
The proof is given in Appendix B. Note that this property is preserved even for the Newton–Raphson method.
3.2 Robustness and the variational representation
The robustness is attributed to the variation representation of the spectral Rényi divergence. Hereafter we additionally assume that $\widetilde{I}_{n}$ converges to $\mathbb{E}[I_{n}]$ in $L_{2}$.
$=\frac{1}{n}\frac{\alpha S_{\theta}(\omega^{*})}{\alpha S_{\theta}(\omega^{*})+(1-\alpha)\widetilde{I}_{n}(\omega^{*})}\frac{z}{(1-\alpha)z+\alpha S_{\theta}(\omega^{*})+(1-\alpha)\widetilde{I}_{n}(\omega^{*})}-\frac{1}{n}\log(1+\widetilde{I}_{n}^{-1}(\omega^{*})z)$
$\quad\quad+\frac{1}{n}\log(1+(1-\alpha)z\{\alpha S_{\theta}(\omega^{*})+(1-\alpha)\widetilde{I}_{n}(\omega^{*})\}^{-1})$
$=o_{P}(1)$
@@ 4 Simulation studies
  • the minimum spectral Rényi divergence estimators with $\alpha\in\{0.5,0.75,0.9\}$ using the smoothed periodogram ($I^{S}_{n}$);
@@ 4.1 The Brune spectral model with attenuation
Consider the Brune spectral model [2] with attenuation [9], where the spectral density is given by
$S^{\mathrm{BA}}_{\theta}(\omega):=\left[\frac{\sigma^{2}}{\{1+(\omega/\omega_{c})^{2}\}^{2}}\right]\exp(-\omega/Q),\;\;\theta=(\sigma,\omega_{c},Q).$
The Brune spectral model $\sigma^{2}/\{1+(\omega/\omega_{c})^{2}\}^{2}$ is the spectral density model of Matérn process [11] with the spectral decay rate fixed to 2 and often used in the earthquake source estimation [3, 20]. The attenuation factor $\exp(-\omega/Q)$ represents anelasticity of the medium through which the seismic wave propagates.
@@
Table 1: Mean values of biases with standard deviations without any trend. Values closest to zero are underlined.
Initial value $\widehat{\sigma}-\sigma^{*}$ $\widehat{\omega_{c}}-\omega_{c}^{*}$ $\widehat{Q}-Q^{*}$
Rényi ($\alpha=0.50$) $\theta^{(0),1}$0.03 ($\pm$0.05) -0.12 ($\pm$0.07) 0.30 ($\pm$0.15)
Rényi ($\alpha=0.75$) $\theta^{(0),1}$ 0.01 ($\pm$0.06) -0.17 ($\pm$0.06) 0.54 ($\pm$0.22)
Rényi ($\alpha=0.90$) $\theta^{(0),1}$ 0.01 ($\pm$0.06) -0.20 ($\pm$0.06) 0.90 ($\pm$0.33)
Itakura–Saito with $I_{n}$ $\theta^{(0),1}$ 42.6 ($\pm$2.66) 893 ($\pm$50.5) 48.2 ($\pm$2.99)
Itakura–Saito with $I_{n}^{S}$ $\theta^{(0),1}$ 3372 ($\pm$5318) 67386 ($\pm$106260) 5259 ($\pm$8327)
Rényi ($\alpha=0.50$) $\theta^{(0),2}$0.03 ($\pm$0.05) -0.10 ($\pm$0.07) 0.23 ($\pm$0.15)
Rényi ($\alpha=0.75$) $\theta^{(0),2}$ 0.002 ($\pm$0.06) -0.17 ($\pm$0.07) 0.50 ($\pm$0.21)
Rényi ($\alpha=0.90$) $\theta^{(0),2}$ 0.01 ($\pm$0.06) -0.23 ($\pm$0.05) 1.05 ($\pm$0.34)
Itakura–Saito with $I_{n}$ $\theta^{(0),2}$ 0.06 ($\pm$0.10) -0.04 ($\pm$0.13) 0.06 ($\pm$0.17)
Itakura–Saito with $I_{n}^{S}$ $\theta^{(0),2}$0.29 ($\pm$0.30) 1.95 ($\pm$1.89) 1.86 ($\pm$0.98)
Rényi ($\alpha=0.50$) $\theta^{(0),3}$0.10 ($\pm$0.04) 0.51 ($\pm$0.13) -0.30 ($\pm$0.06)
Rényi ($\alpha=0.75$) $\theta^{(0),3}$0.05 ($\pm$0.06) 0.08 ($\pm$0.19) 0.05 ($\pm$0.20)
Rényi ($\alpha=0.90$) $\theta^{(0),3}$0.007 ($\pm$0.06) -0.15 ($\pm$0.13) 0.72 ($\pm$0.35)
Itakura–Saito with $I_{n}$ $\theta^{(0),3}$ 0.10 ($\pm$0.07) 0.42 ($\pm$0.26) -0.28 ($\pm$0.13)
Itakura–Saito with $I_{n}^{S}$ $\theta^{(0),3}$0.29 ($\pm$0.15) 2.09 ($\pm$1.68) 1.50 ($\pm$0.72)
@@
Table 2: Mean values of biases with standard deviations with the trigonometric trends. Values closest to zero are underlined.
Initial value $\widehat{\sigma}-\sigma^{*}$ $\widehat{\omega_{c}}-\omega_{c}^{*}$ $\widehat{Q}-Q^{*}$
Rényi ($\alpha=0.50$) $\theta^{(0),1}$ 0.22 ($\pm$0.07) -0.22 ($\pm$0.05) 0.27 ($\pm$0.15)
Rényi ($\alpha=0.75$) $\theta^{(0),1}$ 0.53 ($\pm$0.08) -0.34 ($\pm$0.04) 0.47 ($\pm$0.19)
Rényi ($\alpha=0.90$) $\theta^{(0),1}$ 1.45 ($\pm$0.10) -0.48 ($\pm$0.03) 0.74 ($\pm$0.31)
Itakura–Saito with $I_{n}$ $\theta^{(0),1}$ 75.8 ($\pm$2.85) 1524 ($\pm$55.86) 60.0 ($\pm$3.04)
Itakura–Saito with $I_{n}^{S}$ $\theta^{(0),1}$ 3405 ($\pm$5318) 68019 ($\pm$106263) 5271 ($\pm$8327)
Rényi ($\alpha=0.50$) $\theta^{(0),2}$ 0.21 ($\pm$0.07) -0.20 ($\pm$0.06) 0.23 ($\pm$0.14)
Rényi ($\alpha=0.75$) $\theta^{(0),2}$ 0.53 ($\pm$0.08) -0.33 ($\pm$0.04) 0.45 ($\pm$0.19)
Rényi ($\alpha=0.90$) $\theta^{(0),2}$ 1.48 ($\pm$0.10) -0.49 ($\pm$0.03) 0.79 ($\pm$0.33)
Itakura–Saito with $I_{n}$ $\theta^{(0),2}$ 3.92 ($\pm$0.07) -0.35 ($\pm$0.02) -0.40 ($\pm$0.02)
Itakura–Saito with $I_{n}^{S}$ $\theta^{(0),2}$ 2.83 ($\pm$0.97) -0.09 ($\pm$0.97) 2.30 ($\pm$0.98)
Rényi ($\alpha=0.50$) $\theta^{(0),3}$ 0.08 ($\pm$0.06) 0.24 ($\pm$0.15) -0.25 ($\pm$0.10)
Rényi ($\alpha=0.75$) $\theta^{(0),3}$ 0.46 ($\pm$0.08) -0.27 ($\pm$0.07) 0.26 ($\pm$0.16)
Rényi ($\alpha=0.90$) $\theta^{(0),3}$ 1.40 ($\pm$0.09) -0.47 ($\pm$0.03) 0.69 ($\pm$0.29)
Itakura–Saito with $I_{n}$ $\theta^{(0),3}$ 3.90 ($\pm$0.07) -0.34 ($\pm$0.02) -0.40 ($\pm$0.02)
Itakura–Saito with $I_{n}^{S}$ $\theta^{(0),3}$ 2.80 ($\pm$0.95) -0.13 ($\pm$0.59) 2.01 ($\pm$0.86)
@@
with $z_{1}=z_{2}=100$, $A_{1}=\pi/4$ and $A_{2}=\pi/8$. We report the results based on different values of $z_{1}=z_{2}$ in Appendix C. We use three different initial values: $\theta^{(0),1}=(1,0.1,1)$, $\theta^{(0),2}=(1,1,1)$, and $\theta^{(0),3}=(1,2,1)$. Setting $\sigma^{*}$ and $Q^{*}$ as initial values does not impact on the behavior of the minimum spectral Rényi divergence estimates; we report the results based on different initial values $\theta^{(0),4}=(0.1,0.1,0.1)$ and $\theta^{(0),5}=(2,2,2)$ in Appendix C.
Figures 1 and 2 display the estimated spectral densities. In each figure, the following representations are used:

### v3

@@ 3.1 Stable optimization path of the Spectral Rényi divergence minimization
We begin with showing that the optimization for the minimum spectral Rényi divergence estimator is robust against to outliers in the frequency domain. Consider the gradient descent update with a sequence of learning rates $\gamma_{1},\gamma_{2},\ldots>0$ of $\widehat{\theta}_{\alpha}$: for the $(m+1)$-th step,
with
$\mathcal{G}_{\alpha}(\theta\,;\,\widetilde{I}_{n})$ $=-\nabla_{\theta}D^{(n)}_{\alpha}[\widetilde{I}_{n}\,:\,S_{\theta}]$
$=\frac{\alpha}{(1-\alpha)n}\sum_{\omega\in\Omega_{n}}\left[1-\frac{S_{\theta}(\omega)}{\alpha S_{\theta}(\omega)+(1-\alpha)\widetilde{I}_{n}(\omega)}\right]\nabla_{\theta}\log S_{\theta}(\omega).$ (7)
For theoretical results, we make the following assumptions.
Assumption 1.
We assume the following:
  1. (1)
    The spectral density $S_{\theta}(\omega)$ is differentiable with respect to $\theta$ and we have
    $\sup_{\theta\in\Theta,\,\omega\in[-\pi,\pi]}\max\{\|\nabla_{\theta}S_{\theta}(\omega)\|,\,\|\nabla_{\theta}\log S_{\theta}(\omega)\|,\,\|(1/S_{\theta}(\omega))\nabla_{\theta}\log S_{\theta}(\omega)\|\}$
    $\qquad\qquad\qquad=:U_{1}<\infty;$
  2. (2)
    The spectral density $S_{\theta}(\omega)$ is twice differentiable with respect to $\theta$ and we have
    $\limsup_{n\to\infty}\left(\sup_{\theta\in\Theta}\frac{1}{n}\sum_{\omega\in\Omega_{n}}\|\nabla^{2}_{\theta}\log S_{\theta}(\omega)\|_{\mathrm{op}}\right)=:U_{2}<\infty,$
    where let $\|\cdot\|_{\mathrm{op}}$ denote the operator norm.
Assumption 2.
For any $z\geq 0$, $\omega^{*}\in\Omega_{n}$, and $m\in\mathbb{N}$, the $m$-th gradient decent update $\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}^{z,\omega^{*}}]$ lies in $\Theta$.
Under Assumptions 1 and 2, the subsequent theorem indicates that for any value of $\alpha\in(0,1)$, the path of the gradient descent with a prefixed step size for $\widehat{\theta}_{\alpha}$ is robust against the presence of outliers in the frequency domain.
Theorem 3 (Stability of optimization path with a fixed learning rate sequence).
Fix arbitrary step size sequence $\{\gamma_{k}>0\,:\,k=1,2,\ldots\}$. Under Assumptions 1 and 2, for any $m\in\mathbb{N}$, we have
$\lim_{n\to\infty}\sup_{z\geq 0,\omega^{*}\in\Omega_{n}}\left\|\widehat{\theta}_{\alpha}^{(m)}[\widetilde{I}^{z,\omega^{*}}_{n}]-\widehat{\theta}_{\alpha}^{(m)}[\widetilde{I}_{n}]\right\|=0\text{ almost surely}.$ (8)
The proof is given in Appendix B.
Remark 1 (Constant step size).
Assumption 1 ensures that we can obtain an upper bound of the smoothness constant of $D^{(n)}_{\alpha}[\widetilde{I}_{n}^{z,\omega^{*}}:S_{\theta}]$ with respect to $\theta$ uniformly in $z\geq 0$: for any $z\geq 0$, we have
$\sup_{\theta\neq\theta^{\prime}}\frac{\|\nabla_{\theta}D^{(n)}_{\alpha}[\widetilde{I}^{z,\omega^{*}}_{n}\,:\,S_{\theta}]-\nabla_{\theta}D^{(n)}_{\alpha}[\widetilde{I}^{z,\omega^{*}}_{n}\,:\,S_{\theta^{\prime}}]\|}{\|\theta-\theta^{\prime}\|}$
$\leq\sup_{z\geq 0}\sup_{\theta\in\Theta}\|\nabla_{\theta}^{2}D_{\alpha}^{(n)}[\widetilde{I}_{n}^{z,\omega^{*}}:S_{\theta}]\|_{\mathrm{op}}$
$\leq L:=\frac{U_{1}^{2}+U_{2}}{1-\alpha}<\infty,$ (9)
where we use the following expression of the Hessian of $D_{\alpha}^{(n)}[\widetilde{I}_{n}:S_{\theta}]$:
$\left(-\nabla_{\theta}^{2}D^{(n)}_{\alpha}[\widetilde{I}_{n}:S_{\theta]}]\right)$
$=-\frac{1}{(1-\alpha)n}\sum_{\omega\in\Omega_{n}}\frac{\alpha S_{\theta}(\omega)}{\alpha S_{\theta}(\omega)+(1-\alpha)\widetilde{I}_{n}(\omega)}$
$\qquad\qquad\qquad\qquad\frac{(1-\alpha)\widetilde{I}_{n}(\omega)}{\alpha S_{\theta}(\omega)+(1-\alpha)\widetilde{I}_{n}(\omega)}\left(\nabla_{\theta}\log S_{\theta}(\omega)\right)\left(\nabla_{\theta}\log S_{\theta}(\omega)\right)^{\top}$
$\quad+\frac{\alpha}{(1-\alpha)n}\sum_{\omega\in\Omega_{n}}\left\{1-\frac{S_{\theta}(\omega)}{\alpha S_{\theta}(\omega)+(1-\alpha)\widetilde{I}_{n}(\omega)}\right\}\nabla^{2}_{\theta}\log S_{\theta}(\omega).$
[…]
$\lim_{m\to\infty}\nabla_{\theta=\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}_{n}^{z,\omega^{*}}]}D_{\alpha}^{(n)}[\widetilde{I}_{n}^{z,\omega^{*}}:S_{\theta}]=0.$
If the further assumption called the Kurdyka–Łojasiewicz inequality [18, 16] holds at all stationary points, that is, points satisfying $\nabla_{\theta}D_{\alpha}^{(n)}[\widetilde{I}_{n}^{z,\omega^{*}}:S_{\theta}]=0$ for the function $\theta\mapsto D_{\alpha}^{(n)}[\widetilde{I}_{n}^{z,\omega^{*}}:S_{\theta}]$, the bounded gradient descent update converges to a stationary point; c.f., Theorem $3.2$ of [1]. Theorem 3 implies that the finite update sequences for different $z\geq 0$ become the same and all stationary points for different $z\geq 0$ become the same as $n\to\infty$, which gives an new aspect into the robust spectral analysis and the optimization theory.
Remark 2 (AR(1) model).
Here we check Assumption 1 for a first-order autoregressive model. The model spectral density is
$S_{\sigma,\rho}(\omega)=\frac{1}{2\pi}\frac{\sigma^{2}}{1-2\rho\cos(\omega)+\rho^{2}},\quad\sigma>0\quad\text{and}\quad\rho\in(-1,1).$
We reparametrize this by using
$\theta_{1}=\log\sigma\quad\text{and}\quad\theta_{2}=\log\frac{1+\rho}{1-\rho}$
so as to remove constraints in the optimization. Then, the gradient of $\log S_{\theta}(\omega)$ with respect to $\theta$ is given by
$\nabla_{\theta}\log S_{\theta}(\omega)=\begin{pmatrix}2\\ (1+\rho)(1-\rho)\{\cos(\omega)-\rho\}/\{1-2\rho\cos(\omega)+\rho^{2}\}\end{pmatrix}.$
The Hessian of $\log S_{\theta}(\omega)$ is a bit complex but is bounded as
$\|\nabla^{2}_{\theta}\log S_{\theta}(\omega)\|_{\mathrm{op}}\leq C/(1-|\rho|)^{4}$
for some absolute positive constant $C$. So, Assumption 1 holds for any bounded parameter space $\Theta$ of $\theta$.
Figure 1: Comparison of optimization paths. The optimization paths of the gradient descent for the $AR(1)$ model, with respect to $(\log\sigma,2\{1/(1+\exp(-\rho))-1/2\})$, are presented under the spectral Rényi divergence minimization and the Itakura–Saito divergence minimization. The left figure shows the optimization paths (curves with light colors) of the spectral Rényi divergence minimization that start from the initial value (the gray circle) and terminate at points after $5000$ iterations (the inverted triangles), where the colors indicates different values of the contamination $z$. The right figure shows the corresponding results for the Itakura–Saito divergence minimization.
Theorem 3 establishes a connection between robustness in statistics and optimization theory, offering a novel perspective in robust statistics. While detailed numerical experiments are presented in Section 4, we briefly illustrate the implications through a simple simulation study here.
We consider contaminated observations defined as
$X_{t}^{\circ}=X_{t}+\sqrt{z}\sin(t\pi/2),\quad t=1,\ldots,1000$
with $X_{t}$ following from AR(1) model having $(\sigma,\rho)$. We perform 5000 iterations of gradient descent with a step size of 0.01 with respect to
$(\log\sigma,2\{1/(1+\exp(-\rho))-1/2\}).$
The left panel of Figure 1 illustrates the optimization paths and endpoints based on $D_{\alpha=0.5}[I_{1000}:S_{\theta}]$, starting from the initial value (gray circle). Different colors represent different contamination levels $z\in\{0,10^{2},10^{4}\}$. The figure highlights the robustness of the optimization paths under varying contamination levels. In contrast, the right figure of Figure 1 shows the optimization paths and the stopping points based on the Itakura–Saito divergence $D_{1}[I_{1000}\,:\,S_{\theta}]$ starting from the initial value (the gray circle). This figure demonstrates the non-robustness of the optimization paths of the Itakura–Saito divergence minimization under contamination. The instability observed here will be analyzed further in the subsequent subsection.
Practically, gradient descent updates commonly use line search methods to optimize the step size dynamically. To manage this situation, we present a guiding theorem for gradient descent updates with line search. In this paper, we employ the Armijo condition for this purpose. Consider an objective function $\mathcal{L}(\theta)$ to be minimized and let $\theta^{(m)}$ be the $m$-th gradient descent update for each $m\in\mathbb{N}$. The Armijo condition with $c\in(0,1)$ selects the step size $\gamma$ in the $m$-th update $\theta^{(m)}$ so as to satisfy
$\mathcal{L}\left(\theta^{(m-1)}-\gamma\nabla_{\theta=\theta^{(m-1)}}\mathcal{L}(\theta)\right)\leq\mathcal{L}\left(\theta^{(m-1)}\right)-c\gamma\left\|\nabla_{\theta=\theta^{(m-1)}}\mathcal{L}(\theta)\right\|^{2},$
which ensures the monotone decrease of the objective function along with the gradient descent update. We denote by $\mathcal{A}[c,\theta,\mathcal{L}]$ the Armijo condition with $c\in(0,1)$ for an objective function $\mathcal{L}(\theta)$ in the update from $\theta$.
Proposition 2 (Stability of line search).
Fix $c\in(0,1)$, $\kappa>0$, and $\underline{\gamma}>0$. Fix also the point $\theta\in\Theta$ such that
$\left\|\nabla_{\theta}D_{\alpha}^{(n)}[\widetilde{I}_{n}:S_{\theta}]\right\|>\kappa.$
[…]
The proof is given in Appendix C. It employs the descent lemma (c.f., [2]) together with the comparison of gradients.
Remark 3 (Stability of optimization path with the Armijo condition).
Let $\{\gamma_{m}[\widetilde{I}_{n}]:m=1,2,\ldots\}$ and $\{\gamma_{m}[\widetilde{I}^{z,\omega^{*}}_{n}]:m=1,2,\ldots\}$ the step size sequences for $D_{\alpha}^{(n)}[\widetilde{I}_{n}:S_{\theta}]$ and $D_{\alpha}^{(n)}[\widetilde{I}_{n}^{z,\omega^{*}}:S_{\theta}]$, respectively. From Proposition 2, we can assume
$\sup_{z\geq 0,\omega^{*}\in\Omega_{n}}|\gamma_{m}[\widetilde{I}_{n}]-\gamma_{m}[\widetilde{I}_{n}^{z,\omega^{*}}]|\to 0\quad\text{as }\quad n\to\infty.$
Under this additional assumption and if we have
$\inf_{k=0,\ldots,m-1}\|\nabla_{\theta=\widehat{\theta}_{\alpha}^{(k)}[\widetilde{I}_{n}]}D_{\alpha}^{(n)}[\widetilde{I}_{n}:S_{\theta}]\|>0,$
we can have for any $m\in\mathbb{N}$,
$\lim_{n\to\infty}\sup_{z\geq 0,\omega^{*}\in\Omega_{n}}\left\|\widehat{\theta}_{\alpha}^{(m)}[\widetilde{I}^{z,\omega^{*}}_{n}]-\widehat{\theta}_{\alpha}^{(m)}[\widetilde{I}_{n}]\right\|=0\text{ almost surely},$
even when we utilize the line search.
3.2 Instable optimization path of the Itakura–Saito divergence minimization
We next show that the optimization for the minimum Itakura–Saito divergence estimator is sensitive to outliers in the frequency domain. Consider the gradient descent update with a sequence of learning rates $\gamma_{1},\gamma_{2},\ldots>0$ of $\widehat{\theta}_{1}$: for the $(m+1)$-th step,
@@ 3.2 Instable optimization path of the Itakura–Saito divergence minimization
The proof is given in Appendix D.
@@ 4 Simulation studies
  • the minimum spectral Rényi divergence estimators with
    $\alpha\in\{0.5,0.75,0.9\}$
    using the smoothed periodogram ($I^{S}_{n}$);
@@ 4.1 The Brune spectral model with attenuation
Table 1: Mean values of biases with standard deviations without any trend. Values closest to zero are underlined. Rényi is abbreviated as R; Itakura–Saito is abbreviated as IS; Initial value is abbreviated as Init.
Init $\widehat{\sigma}-\sigma^{*}$ $\widehat{\omega_{c}}-\omega_{c}^{*}$ $\widehat{Q}-Q^{*}$
R ($\alpha=0.50$) $\theta^{(0),1}$0.03 ($\pm$0.05) -0.12 ($\pm$0.07) 0.30 ($\pm$0.15)
R ($\alpha=0.75$) $\theta^{(0),1}$ 0.01 ($\pm$0.06) -0.17 ($\pm$0.06) 0.54 ($\pm$0.22)
R ($\alpha=0.90$) $\theta^{(0),1}$ 0.01 ($\pm$0.06) -0.20 ($\pm$0.06) 0.90 ($\pm$0.33)
IS with $I_{n}$ $\theta^{(0),1}$ 42.6 ($\pm$2.66) 893 ($\pm$50.5) 48.2 ($\pm$2.99)
IS with $I_{n}^{S}$ $\theta^{(0),1}$ 3372 ($\pm$5318) 67386 ($\pm$106260) 5259 ($\pm$8327)
R ($\alpha=0.50$) $\theta^{(0),2}$0.03 ($\pm$0.05) -0.10 ($\pm$0.07) 0.23 ($\pm$0.15)
R ($\alpha=0.75$) $\theta^{(0),2}$ 0.002 ($\pm$0.06) -0.17 ($\pm$0.07) 0.50 ($\pm$0.21)
R ($\alpha=0.90$) $\theta^{(0),2}$ 0.01 ($\pm$0.06) -0.23 ($\pm$0.05) 1.05 ($\pm$0.34)
IS with $I_{n}$ $\theta^{(0),2}$ 0.06 ($\pm$0.10) -0.04 ($\pm$0.13) 0.06 ($\pm$0.17)
IS with $I_{n}^{S}$ $\theta^{(0),2}$0.29 ($\pm$0.30) 1.95 ($\pm$1.89) 1.86 ($\pm$0.98)
R ($\alpha=0.50$) $\theta^{(0),3}$0.10 ($\pm$0.04) 0.51 ($\pm$0.13) -0.30 ($\pm$0.06)
R ($\alpha=0.75$) $\theta^{(0),3}$0.05 ($\pm$0.06) 0.08 ($\pm$0.19) 0.05 ($\pm$0.20)
R ($\alpha=0.90$) $\theta^{(0),3}$0.007 ($\pm$0.06) -0.15 ($\pm$0.13) 0.72 ($\pm$0.35)
IS with $I_{n}$ $\theta^{(0),3}$ 0.10 ($\pm$0.07) 0.42 ($\pm$0.26) -0.28 ($\pm$0.13)
IS with $I_{n}^{S}$ $\theta^{(0),3}$0.29 ($\pm$0.15) 2.09 ($\pm$1.68) 1.50 ($\pm$0.72)
@@
Table 2: Mean values of biases with standard deviations with the trigonometric trends. Values closest to zero are underlined. Rényi is abbreviated as R; Itakura–Saito is abbreviated as IS; Initial value is abbreviated as Init.
Init $\widehat{\sigma}-\sigma^{*}$ $\widehat{\omega_{c}}-\omega_{c}^{*}$ $\widehat{Q}-Q^{*}$
R ($\alpha=0.50$) $\theta^{(0),1}$ 0.22 ($\pm$0.07) -0.22 ($\pm$0.05) 0.27 ($\pm$0.15)
R ($\alpha=0.75$) $\theta^{(0),1}$ 0.53 ($\pm$0.08) -0.34 ($\pm$0.04) 0.47 ($\pm$0.19)
R ($\alpha=0.90$) $\theta^{(0),1}$ 1.45 ($\pm$0.10) -0.48 ($\pm$0.03) 0.74 ($\pm$0.31)
IS with $I_{n}$ $\theta^{(0),1}$ 75.8 ($\pm$2.85) 1524 ($\pm$55.86) 60.0 ($\pm$3.04)
IS with $I_{n}^{S}$ $\theta^{(0),1}$ 3405 ($\pm$5318) 68019 ($\pm$106263) 5271 ($\pm$8327)
R ($\alpha=0.50$) $\theta^{(0),2}$ 0.21 ($\pm$0.07) -0.20 ($\pm$0.06) 0.23 ($\pm$0.14)
R ($\alpha=0.75$) $\theta^{(0),2}$ 0.53 ($\pm$0.08) -0.33 ($\pm$0.04) 0.45 ($\pm$0.19)
R ($\alpha=0.90$) $\theta^{(0),2}$ 1.48 ($\pm$0.10) -0.49 ($\pm$0.03) 0.79 ($\pm$0.33)
IS with $I_{n}$ $\theta^{(0),2}$ 3.92 ($\pm$0.07) -0.35 ($\pm$0.02) -0.40 ($\pm$0.02)
IS with $I_{n}^{S}$ $\theta^{(0),2}$ 2.83 ($\pm$0.97) -0.09 ($\pm$0.97) 2.30 ($\pm$0.98)
R ($\alpha=0.50$) $\theta^{(0),3}$ 0.08 ($\pm$0.06) 0.24 ($\pm$0.15) -0.25 ($\pm$0.10)
R ($\alpha=0.75$) $\theta^{(0),3}$ 0.46 ($\pm$0.08) -0.27 ($\pm$0.07) 0.26 ($\pm$0.16)
R ($\alpha=0.90$) $\theta^{(0),3}$ 1.40 ($\pm$0.09) -0.47 ($\pm$0.03) 0.69 ($\pm$0.29)
IS with $I_{n}$ $\theta^{(0),3}$ 3.90 ($\pm$0.07) -0.34 ($\pm$0.02) -0.40 ($\pm$0.02)
IS with $I_{n}^{S}$ $\theta^{(0),3}$ 2.80 ($\pm$0.95) -0.13 ($\pm$0.59) 2.01 ($\pm$0.86)
Consider the Brune spectral model [4] with attenuation [14] for velocity waveforms:
$S^{\mathrm{BA}}_{\theta}(\omega):=\omega^{2}\left[\frac{\sigma^{2}}{\{1+(\omega/\omega_{c})^{2}\}^{2}}\right]\exp(-\omega/Q),\;\;\theta=(\sigma,\omega_{c},Q).$
The Brune spectral model $\sigma^{2}/\{1+(\omega/\omega_{c})^{2}\}^{2}$ for displacement waveforms is the spectral density model of Matérn process [21] with the spectral decay rate fixed to 2 and often used in the earthquake source estimation [5, 33]. The attenuation factor $\exp(-\omega/Q)$ represents anelasticity of the medium through which the seismic wave propagates. The multiplicative factor $\omega^{2}$ comes from the time derivative that converts a displacement waveform to a velocity waveform.
@@
with $z_{1}=z_{2}=100$, $A_{1}=\pi/4$ and $A_{2}=\pi/8$. We report the results based on different values of $z_{1}=z_{2}$ in Appendix E. We use three different initial values: $\theta^{(0),1}=(1,0.1,1)$, $\theta^{(0),2}=(1,1,1)$, and $\theta^{(0),3}=(1,2,1)$. Setting $\sigma^{*}$ and $Q^{*}$ as initial values does not impact on the behavior of the minimum spectral Rényi divergence estimates; we report the results based on different initial values $\theta^{(0),4}=(0.1,0.1,0.1)$ and $\theta^{(0),5}=(2,2,2)$ in Appendix E.
Figures 2 and 3 display the estimated spectral densities. In each figure, the following representations are used: